Question 10 (1 point) Heinlein Corporation is financed with 0.5 percent debt and the rest equity. It has a leveraged beta of 1.1 and is subject to a 0.1 corporate tax rate. What is Heinlein's unleveraged beta? Your Answer:

Question 10 options: Answer Question 11 (1 point) Empress Corporation has a capital budget of 140 and and a Net Income of 360. If Empress's target equity fraction is 0.4, what should Empress's dividend be under the residual dividend model?
Answer #1

10). Unlevered Beta = Levered Beta / [1 + {(D/E) * (1 - t)}]

= 1.1 / [1 + {(0.5/0.5) * (1 - 0.1)}]

= 1.1 / [1 + 0.9] = 1.1 / 1.9 = 0.58

11). Equity Requirement = Capital Budget * % of Equity = 140 * 0.4 = 56

Dividends = Net Income - Equity Requirement = 360 - 56 = 304